OpenStudy (anonymous):

Find the fundamental period of f when f(x) = 3 cos(5x)

OpenStudy (anonymous):

period of \[\cos(bx)\] is \[\frac{2\pi}{b}\]

OpenStudy (anonymous):

i assume "fundamental period" is a synonym for "period'

OpenStudy (anonymous):

I wasn't sure , i know how to find the period but i dont know if it is the same thing

OpenStudy (anonymous):

a google search tells me it is the smallest number p for which \[f(x)=f(x+p)\] so yes it is the same as in common usage

OpenStudy (anonymous):

ooh ok